The sum of two numbers is 40. When the greater number is divided by the smaller, the quotient is 4 and the remainder is 5. Find the numbers.
Answer by themathtutor2009(81)

x+y=40....let x > y

x/y = 4+5/y

x = 4y+5


x+y=40
4y+5+y=40
5y+5=40
5y=35
y=7
x=4y+5=4(7)+5=28+5=33

Answers: numbers are 33 and 7